OpenStudy (anonymous):

A car wheel has a 16-inch radius. Through what angle (to the nearest tenth of a degree) does the wheel turn when the car rolls forward 2 ft?

OpenStudy (inkyvoyd):

(2*12)/(16*2pi)*360

OpenStudy (anonymous):

Is there a certain formula you used to arrive at that equation? I know it has to do with arc length, but I get stuck after that.

OpenStudy (dumbcow):

oops i didn't see the units, 2ft = 24 in 24 = 16A

OpenStudy (dumbcow):

--> A = 24/16 = 3/2 conveet to degrees by multiplying by (180/pi) 3/2 * 180/pi = 85.94 degrees

OpenStudy (anonymous):

Ok, slow on the uptake, sorry, but if "s" equals the arc length, "r" equals the radius (16 inches in this case), and theta equals the central angle, where does the two feet fit into the problem here? s=r*(theta). How can I fill in the blanks of this formula?

OpenStudy (dumbcow):

2 ft or 24 in is the arc length

OpenStudy (dumbcow):